# Gravitational Positivity Bounds on Higgs-Portal Light Dark Matter

### v2

@@ Abstract
Gravitational positivity bounds are constraints on a renormalizable theory in the presence of a massless graviton, under the assumption that the gravitational theory is ultraviolet-completed by a perturbative string theory. We derive these bounds for the Higgs-portal scalar dark matter model using the forward scattering process $\phi\phi\to\phi\phi$. We find that, in the absence of a dark matter self-coupling, new physics beyond the Higgs-portal dark matter interaction must appear below an energy scale of $10^{10}$ GeV if the dark matter mass is smaller than the Higgs boson mass. The presence of a dark matter self-coupling alters this situation. A hierarchy between the dark matter four-point self-coupling $\lambda_{\phi}$ and a tiny Higgs-portal coupling $\lambda_{h\phi}$ is required to raise the energy scale at which the new physics appears. If $\lambda_{\phi}/\lambda_{h\phi}=10^{12}$, the dark matter model can remain valid up to the grand unified theory (GUT) scale or the typical string scale. In this case, the relic abundance of dark matter in the Universe can be reproduced via the dark freeze-out scenario. A parameter set with $\lambda_{\phi}\sim\mathcal{O}(1)$, $\lambda_{h\phi}\sim 10^{-12}$, and a sub-GeV dark matter mass can accommodate the GUT-scale $\Lambda$ within the Higgs-portal light dark matter framework.
@@ 1 Introduction
In this paper, we derive gravitational positivity bounds for the renormalizable Higgs-portal scalar DM interactions and discuss their phenomenological implications. This paper is organized as follows. In Section 2, we introduce the Lagrangian of the model. In Section 3, we review the derivation of gravitational positivity bounds for renormalizable interactions and present the bounds obtained for the Higgs-portal scalar DM model. In Section 4, we discuss the phenomenological implications of these bounds for the Higgs-portal real scalar DM model. Finally, in Section 5, we summarize our findings and outline directions for future work.
@@ 2 Higgs-Portal Scalar Dark Matter
3 Gravitational Positivity Bounds and Application for Higgs-Portal Scalar DM
@@ 3.1 Pedagogical Review
We review the derivation of gravitational positivity bounds using the tree level string-state contributions to the $\phi\phi\to\phi\phi$ scattering process. Although we explicitly consider the $\phi\phi\to\phi\phi$ process as an example, the discussion is generic. We introduce two scales $\Lambda<\Lambda_{\text{QG}}$ above the electroweak scale, the new physics scale $\Lambda$ and the quantum gravity scale $\Lambda_{\text{QG}}$, and assume the following:
@@
Here, we include the $u$-channel contributions in the LHS of Eq. (11) via $u=4m_{\phi}^{2}-s-t\sim 4m_{\phi}^{2}-s$, which are absorbed into $\tilde{a}_{2}(t)$ and $g_{2}(t)$, respectively. The right-hand side of Eq. (11) can be divided into three energy regions:
@@
In the forward limit $t\to-0$, the first term on the right-hand side of Eq. (14) diverges to -$\infty$ since $C>0$. However, this divergence is canceled by the contributions of tree-level Regge states in the first integral, corresponding to energies $E>\Lambda_{\mathrm{QG}}$. Above the string or quantum gravity scale, $\Lambda_{\mathrm{QG}}=\alpha^{\prime-1/2}$, a tower of massive higher spin states (a Regge tower) appear, as assumed in assumption (i). The scattering amplitude receives contributions from these Regge states [56]. Here, the scale $\alpha^{\prime}$ corresponds to the usual $\alpha^{\prime}$ parameter in string theory. The associated divergence is almost canceled by summing the contributions of the Regge states, leaving at most a possible remnant of order $\mathcal{O}(\alpha^{\prime}/M^{2}_{\mathrm{pl}})=\mathcal{O}(1/(\Lambda_{\mathrm{QG}}^{2}M^{2}_{\mathrm{pl}}))$, as shown by Tokuda et al. [56] using string theory. (See also Refs. [13, 34].)
@@
Here, we take the forward limit $t\to-0$. Note that the right-hand side of Eq. (15) is manifestly positive due to unitarity, assumption (ii), since unitarity leads to the optical theorem for forward 2-to-2 elastic scattering:
@@
Here, $B_{\text{grav}}$ represents the subtraction of the $s^{2}$ dependence of the amplitude, so its mass dimension is 4-4. Non-gravitational part, $B_{\text{non-grav}}$, likewise has mass dimension 4-4, and it involves the subtraction of a positive contribution from the integrand.
@@
Here, $\tilde{a}_{2}(t\to-0)$ corresponds to a combination of the Wilson coefficients of dimension-8 operators (for example, see [57]). Eq. (22) shows that these dimension-8 contributions on the LHS arise from heavy states appearing at or above the energy scale $\Lambda$. By unitarity of the new physics sector, these states provide positive contributions to the RHS via $\mathrm{Im}\,\mathcal{M}$. This is the standard argument for the original positivity bounds [3].
@@ 3.2 Gravitational Positivity Bounds with $\phi\phi\to\phi\phi$
We now compute $B(\Lambda)$ for the $\phi\phi\to\phi\phi$ scattering process at the one-loop level, following Sec. 3.1. The one-loop amplitudes are calculated using FeynRules [25, 7] and the Mathematica packages FeynArts [33], FeynCalc [45, 52, 53, 54], and Package-X [50].
@@
$B^{\phi\phi\to\phi\phi}_{\text{grav}}$ $=B^{\phi\phi\to\phi\phi}_{\text{grav(a)}}+B^{\phi\phi\to\phi\phi}_{\text{grav(b)}}+B^{\phi\phi\to\phi\phi}_{\text{grav(c)}}+B^{\phi\phi\to\phi\phi}_{\text{grav(d)}}=-\frac{\lambda^{2}_{h\phi}v^{2}}{24\pi^{2}\overline{M}^{2}_{\text{pl}}m^{4}_{h}}f(m_{\phi}/m_{h}),$ (24)
@@
where $\overline{M}_{\text{pl}}=M_{\text{pl}}/\sqrt{8\pi}\sim 2.4\times 10^{18}~\mathrm{GeV}$ is the reduced Planck mass, and $m_{h}$ is the Higgs boson mass. Explicit expressions for each contribution, i.e., $B^{\phi\phi\to\phi\phi}_{\text{grav(a/b/c/d)}}$, are provided in Appendix B.
@@
Combining Eqs. (23) and (24), the gravitational positivity bound for the $\phi\phi\to\phi\phi$ process is
$B^{\phi\phi\to\phi\phi}(\Lambda)$ $=B^{\phi\phi\to\phi\phi}_{\text{non-grav}}+B^{\phi\phi\to\phi\phi}_{\text{grav}}\geq 0,$ (26)
$\lambda_{h\phi}^{2}+\lambda_{\phi}^{2}$ $\geq\frac{2\lambda^{2}_{h\phi}\Lambda^{4}v^{2}}{3\overline{M}^{2}_{\text{pl}}m^{4}_{h}}f(m_{\phi}/m_{h}),$ (27)
which explicitly gives
$\Lambda$ $\leq\left(\frac{3}{2}\right)^{1/4}\left(\frac{\overline{M}_{\text{pl}}}{v}\right)^{1/2}\left(1+\frac{\lambda^{2}_{\phi}}{\lambda^{2}_{h\phi}}\right)^{1/4}f^{-1/4}(m_{\phi}/m_{h})m_{h}\,,\quad\text{or}\quad\lambda_{h\phi}=0.$ (28)
@@
$\Lambda$ $\lesssim 10^{10}~\mathrm{GeV}.$ (30)
@@
From Eq. (30), new physics is expected to appear below approximately $10^{10}$ GeV when the Higgs-portal scalar DM has no self-interaction. Here, we assume that $f^{-1/4}(m_{\phi}/m_{h})$ is of order $\mathcal{O}(1)$. Note that $f(x)\to 1/3$ as $x\to 0$, corresponding to $f^{-1/4}(x)\sim 1.3$. Therefore, we focus on the light DM region. The function $f^{-1/4}(x)$, shown in the right panel of Figure 4, takes values of order $\mathcal{O}(1)$ when $m_{\phi}<m_{h}$, supporting our focus on this DM mass region. As the figure suggests, heavier DM masses may allow $\Lambda$ to reach the GUT scale. Indeed, this occurs if the DM mass exceeds approximately $10^{10}$ GeV. However, the light DM region with $m_{\phi}<m_{h}$ is more interesting to investigate, as it is more severely constrained by the gravitational positivity bounds.
@@
Figure 5: Parameter space for $\lambda_{\phi}/\lambda_{h\phi}$ versus $\Lambda$ for light DM with $m_{\phi}<m_{h}$, satisfying the gravitational positivity bound from the $\phi\phi\to\phi\phi$ process. The green region indicates where the bound in Eq. (28) is satisfied.
We illustrated the parameter space of $\lambda_{\phi}/\lambda_{h\phi}$ versus $\Lambda$ in Figure 5 for non-zero $\lambda_{\phi}$. The gravitational positivity bounds in Eq. (28) are satisfied in the green region. As expected from Eq. (28), comparing the cases $\lambda_{\phi}=0$ and $\lambda_{\phi}/\lambda_{h\phi}=100$ in Figure 5, the upper bound on the new physics scale $\Lambda$ increases by approximately one order of magnitude. If there is a large hierarchy between the two couplings, such that the Higgs-portal coupling is small (i.e., $\lambda_{\phi}/\lambda_{h\phi}=10^{2n}$ with $n=1,2,\dots$), the upper bound on the new physics scale is approximately given by
$\Lambda$ $\lesssim 10^{10+n}~\mathrm{GeV},$
$\lambda_{\phi}/\lambda_{h\phi}$ $=10^{2n}\quad(n=1,2,\cdots),$
$\text{for }\,m_{\phi}$ $<m_{h}.$ (31)
4 Phenomenological Implications for Light Dark Matter
Based on the results in Sec. 3, we discuss the implications of the gravitational positivity bounds for the Higgs-portal light real scalar DM model. From Eq. (31), if the new physics scale $\Lambda$ corresponds to the GUT scale or a typical string scale, i.e. $\Lambda\sim 10^{16}$ GeV, the bound implies a large hierarchy between the Higgs-portal coupling and the DM self-coupling:
$\lambda_{\phi}/\lambda_{h\phi}$ $=10^{12}\quad(\text{for $\Lambda\sim$ GUT/typical string scale}),$
$\text{for }m_{\phi}$ $<m_{h}.$ (32)
From Eq. (32) and requiring perturbativity, i.e., $\lambda_{\phi}\lesssim 4\pi$, it follows that $\lambda_{h\phi}\ll 1$. This implies that the Higgs-portal coupling must be extremely small, if it exists at all. When $\lambda_{h\phi}\ll 1$ but non-zero, the Higgs-portal DM freeze-in scenario becomes relevant [19, 41].
We consider two cases for light DM ($m_{\phi}<m_{h}$): $\lambda_{\phi}\lesssim 10^{-2}$ and $\lambda_{\phi}\gtrsim 10^{-2}$. The latter case is necessary to realize the GUT-scale or typical string-scale $\Lambda$.
  • Case 1: ($\lambda_{\phi}\lesssim 10^{-2}$)
    For light DM, i.e., $m_{\phi}<m_{h}$, the process $h\to\phi\phi$ dominates over $hh\to\phi\phi$ [41] in determining the DM relic abundance. For 100 MeV DM [19],
    $\lambda_{\phi}/\lambda_{h\phi}\lesssim 10^{10}.$ (33)
    This corresponds to $\Lambda\lesssim 10^{15}$ GeV according to Eq. (31). Since a smaller DM mass lowers $\Lambda$, this scenario does not lead to a GUT-scale $\Lambda$.
  • Case 2: ($\lambda_{\phi}\gtrsim 10^{-2}$)
    Here, the DM self-interaction plays a significant role in DM production via dark freeze-out [19, 31, 2]. With a non-zero Higgs-portal coupling, the following parameter set can simultaneously realize both the GUT-scale $\Lambda$ (or typical string scale $\Lambda$) and the correct DM relic abundance (see Fig. 5 in [19]):
    $\lambda_{h\phi}$ $\sim 2\times 10^{-12},$
    $\lambda_{\phi}$ $\sim 2,$
    $m_{\phi}$ $\sim 300~\mathrm{MeV}.$ (34)
    Additionally, this parameter set (34) satisfies stringent astrophysical constraints from Milky Way satellite dwarf galaxies, i.e., $\sigma/m<0.2~\mathrm{cm}^{2}/\mathrm{g}$ [9], where $\sigma/m$ is the DM self-interaction cross section divided by its mass.
    The dark freeze-out scenario proceeds as follows [19]. At temperatures below the Higgs mass (but still relatively high), Higgs decays ($h\to\phi\phi$) produce an initial population of DM particles. Once a sufficient number of DM particles is produced, $2\to 4$ DM self-interactions rapidly increase the DM number density. After chemical equilibrium is reached, the DM number density tracks the equilibrium value until the particles become non-relativistic. Eventually, as the temperature $T$ (or dark temperature $T_{D}$) decreases, freeze-out occurs when $m_{\phi}/T\sim 0.3$ (or $m_{\phi}/T_{D}\sim 7$), fixing the DM yield.
In conclusion, the gravitational positivity bound for light DM, i.e., $m_{\phi}<m_{h}$, indicates that realizing the GUT-scale or typical string-scale $\Lambda$ in the Higgs-portal model requires a large hierarchy between an $\mathcal{O}(1)$ DM self-coupling and an extremely small Higgs-portal coupling of $\mathcal{O}(10^{-12})$, together with a sub-GeV DM mass.

### v3

# Gravitational Positivity Bounds on Higgs-Portal Dark Matter

@@ Abstract
Gravitational positivity bounds are constraints on a renormalizable theory in the presence of a massless graviton, under the assumption that the gravitational theory is ultraviolet-completed by a perturbative string theory. We derive these bounds for the Higgs-portal scalar dark matter model using the forward scattering process $\phi\phi\to\phi\phi$. We find that, in the absence of a dark matter self-coupling, new physics beyond the Higgs-portal dark matter interaction must appear below an energy scale of $10^{10}\,\mathrm{GeV}$ if the dark matter mass is smaller than the Higgs boson mass. We further find that, in the presence of both interactions, achieving a cutoff scale at the grand unified theory scale generally requires a dark matter mass of order $10^{10}$$10^{11}\,\mathrm{GeV}$ (or above), with larger values favored when the four-point self-coupling plays a significant role. For such heavy Higgs-portal dark matter, the observed relic abundance of dark matter in the Universe can be successfully reproduced via the freeze-in mechanism with a tiny Higgs-portal coupling, $\lambda_{h\phi}\lesssim 3.5\times 10^{-11}$. The reheating temperature is then constrained to be $T_{\mathrm{reh}}\lesssim 10^{14}\,\mathrm{GeV}$ by the positivity bounds on the dark matter mass.
@@ 1 Introduction
In this paper, we derive gravitational positivity bounds for the renormalizable Higgs-portal scalar DM interactions and discuss their phenomenological implications. The remainder of this paper is organized as follows. In Section 2, we introduce the Lagrangian of the model. In Section 3, we review the derivation of gravitational positivity bounds for renormalizable interactions and present the bounds obtained for the Higgs-portal scalar DM model. In Section 4, we discuss the phenomenological implications of these bounds. Finally, in Section 5, we summarize our findings and outline directions for future work.
@@ 2 Higgs-Portal Scalar Dark Matter
3 Gravitational Positivity Bounds and Application to Higgs-Portal Scalar DM
@@ 3.1 Pedagogical Review
We review the derivation of gravitational positivity bounds using the tree-level string-state contributions to the $\phi\phi\to\phi\phi$ scattering process. Although we explicitly consider the $\phi\phi\to\phi\phi$ process as an example, the discussion is generic. We introduce two scales $\Lambda<\Lambda_{\text{QG}}$ above the electroweak scale, the new physics scale $\Lambda$ and the quantum gravity scale $\Lambda_{\text{QG}}$, and assume the following:
@@
Here we include the $u$-channel contributions in the LHS of Eq. (11) via $u=4m_{\phi}^{2}-s-t\sim 4m_{\phi}^{2}-s$, which are absorbed into $\tilde{a}_{2}(t)$ and $g_{2}(t)$, respectively. The right-hand side (RHS) of Eq. (11) can be divided into three energy regions:
@@
In the forward limit $t\to-0$, the first term on the right-hand side of Eq. (14) diverges to -$\infty$ since $C>0$. However, this divergence is canceled by the contributions of tree-level Regge states in the first integral, corresponding to energies $E>\Lambda_{\mathrm{QG}}$. Above the string or quantum gravity scale, $\Lambda_{\mathrm{QG}}=\alpha^{\prime-1/2}$, a tower of massive higher spin states (a Regge tower) appear, as assumed in assumption (i). The scattering amplitude receives contributions from these Regge states [14]. Here the scale $\alpha^{\prime}$ corresponds to the usual $\alpha^{\prime}$ parameter in string theory. The associated divergence is almost canceled by summing the contributions of the Regge states, leaving at most a possible remnant of order $\mathcal{O}(\alpha^{\prime}/M^{2}_{\mathrm{pl}})=\mathcal{O}(1/(\Lambda_{\mathrm{QG}}^{2}M^{2}_{\mathrm{pl}}))$, as shown by Tokuda et al. [14] using string theory. (See also Refs. [19, 25].)
@@
Here we take the forward limit $t\to-0$. Note that the right-hand side of Eq. (15) is manifestly positive due to unitarity, assumption (ii), since unitarity leads to the optical theorem for forward 2-to-2 elastic scattering:
@@
Here $B_{\text{grav}}$ represents the subtraction of the $s^{2}$ dependence of the amplitude, so its mass dimension is 4-4. Non-gravitational part, $B_{\text{non-grav}}$, likewise has mass dimension 4-4, and it involves the subtraction of a positive contribution from the integrand.
@@
Here $\tilde{a}_{2}(t\to-0)$ corresponds to a combination of the Wilson coefficients of dimension-8 operators (for example, see [42]). Eq. (22) shows that these dimension-8 contributions on the LHS arise from heavy states appearing at or above the energy scale $\Lambda$. By unitarity of the new physics sector, these states provide positive contributions to the RHS via $\mathrm{Im}\,\mathcal{M}$. This is the standard argument for the original positivity bounds [10].
@@ 3.2 Gravitational Positivity Bounds with $\phi\phi\to\phi\phi$
We now compute $B(\Lambda)$ for the $\phi\phi\to\phi\phi$ scattering process at the one-loop level, following the procedure described in Sec. 3.1. The one-loop amplitudes are calculated using FeynRules [43, 44] and the Mathematica packages FeynArts [45], FeynCalc [46, 47, 48, 49], and Package-X [50].
@@
$B^{\text{1-loop}}_{\text{grav}}$ $=B^{\text{1-loop}}_{\text{grav(a)}}+B^{\text{1-loop}}_{\text{grav(b)}}+B^{\text{1-loop}}_{\text{grav(c)}}+B^{\text{1-loop}}_{\text{grav(d)}}=-\frac{\lambda^{2}_{h\phi}v^{2}}{24\pi^{2}\overline{M}^{2}_{\text{pl}}m^{4}_{h}}f(m_{\phi}/m_{h}),$ (24)
@@
where $\overline{M}_{\text{pl}}=M_{\text{pl}}/\sqrt{8\pi}\sim 2.4\times 10^{18}~\mathrm{GeV}$ is the reduced Planck mass, and $m_{h}$ is the Higgs boson mass. Explicit expressions for each contribution, i.e., $B^{\text{1-loop}}_{\text{grav(a/b/c/d)}}$, are provided in Appendix B.
@@
Figure 5: Two-loop contribution to the forward scattering process $\phi\phi\to\phi\phi$, involving only the real scalar DM field $\phi$, corresponding to the leading contribution from the four-point self-coupling in the gravitational positivity bound.
From Eq. (24), we can see that only the Higgs-portal coupling contributes to the gravitational correction $B^{\text{1-loop}}_{\text{grav}}$. The contribution from the four-point self-coupling first appears at the two-loop level and provides the leading contribution, as depicted in Figure 5 and given by [17]
$B^{\text{2-loop}}_{\text{grav}}=-\frac{10-\pi^{2}}{4608\pi^{4}}\frac{\lambda^{2}_{\phi}}{\overline{M}^{2}_{\text{pl}}m^{2}_{\phi}}.$ (26)
Then,
$B^{\phi\phi\to\phi\phi}_{\text{grav}}$ $=B^{\text{1-loop}}_{\text{grav}}+B^{\text{2-loop}}_{\text{grav}}$ (27)
$=-\frac{\lambda^{2}_{h\phi}v^{2}}{24\pi^{2}\overline{M}^{2}_{\text{pl}}m^{4}_{h}}f(m_{\phi}/m_{h})-\frac{10-\pi^{2}}{4608\pi^{4}}\frac{\lambda^{2}_{\phi}}{\overline{M}^{2}_{\text{pl}}m^{2}_{\phi}}.$ (28)
Combining Eqs. (23) and (28), the gravitational positivity bound for the $\phi\phi\to\phi\phi$ process is given by
$B^{\phi\phi\to\phi\phi}(\Lambda)$ $=B^{\phi\phi\to\phi\phi}_{\text{non-grav}}+B^{\phi\phi\to\phi\phi}_{\text{grav}}\geq 0,$
$\frac{\lambda_{h\phi}^{2}+\lambda_{\phi}^{2}}{16\pi^{2}\Lambda^{4}}$ $\geq\frac{\lambda^{2}_{h\phi}v^{2}}{24\pi^{2}\overline{M}^{2}_{\text{pl}}m^{4}_{h}}f(m_{\phi}/m_{h})+\frac{10-\pi^{2}}{4608\pi^{4}}\frac{\lambda^{2}_{\phi}}{\overline{M}^{2}_{\text{pl}}m^{2}_{\phi}},$ (29)
$\lambda_{h\phi}^{2}+\lambda_{\phi}^{2}$ $\geq\frac{2\Lambda^{4}}{3\overline{M}^{2}_{\text{pl}}}\left(\frac{\lambda^{2}_{h\phi}v^{2}}{m^{4}_{h}}f(m_{\phi}/m_{h})+\frac{10-\pi^{2}}{192\pi^{2}}\frac{\lambda^{2}_{\phi}}{m^{2}_{\phi}}\right),$ (30)
which explicitly leads to the upper bound
$\Lambda\leq\left[\frac{3\,\overline{M}_{\mathrm{pl}}^{2}m_{h}^{4}m_{\phi}^{2}}{2}\,\frac{192\pi^{2}\left(\lambda_{h\phi}^{2}+\lambda_{\phi}^{2}\right)}{192\pi^{2}\lambda_{h\phi}^{2}v^{2}m_{\phi}^{2}f\!\left(\frac{m_{\phi}}{m_{h}}\right)+(10-\pi^{2})\lambda_{\phi}^{2}m_{h}^{4}}\right]^{1/4}.$ (31)
@@
$\Lambda$ $\lesssim 10^{10}~\mathrm{GeV}\,\text{ for }\,m_{\phi}<m_{h}.$ (33)
@@
Equation (33) indicates that new physics is expected to appear below an energy scale of approximately $10^{10}$ GeV when the Higgs-portal scalar DM has no self-interaction and a mass satisfying $m_{\phi}<m_{h}$. Here we assume that $f^{-1/4}(m_{\phi}/m_{h})$ is of order $\mathcal{O}(1)$. Note that $f(x)\to 1/3$ as $x\to 0$, which corresponds to $f^{-1/4}(x)\simeq 1.3$. Therefore, this assumption is valid in the light DM regime. The function $f^{-1/4}(x)$, shown in the right panel of Figure 4, indeed takes values of order $\mathcal{O}(1)$ for $m_{\phi}<m_{h}$, supporting the estimate in Eq. (33). As the figure suggests, a heavier DM mass allows the cutoff scale $\Lambda$ to approach the GUT scale, i.e., $10^{16}$ GeV. Indeed, this occurs once the DM mass exceeds approximately $10^{10}$ GeV.
@@
Figure 6: Parameter space of $m_{\phi}/m_{h}$ versus $\Lambda$ satisfying the gravitational positivity bound from the $\phi\phi\to\phi\phi$ process. The left (right) panel corresponds to $\lambda_{\phi}=0$ ($\lambda_{h\phi}=0$). The green region indicates the parameter space where the bound in Eq. (31) is satisfied.
The parameter space of $m_{\phi}/m_{h}$ versus $\Lambda$ is illustrated in Figure 6. The left panel corresponds to the pure Higgs-portal case, $\lambda_{\phi}=0$, whereas the right panel represents the pure DM four-point coupling case, $\lambda_{h\phi}=0$. In both panels, the green shaded region satisfies the gravitational positivity bound given in Eq. (31). As anticipated from Eq. (31), a comparison between the cases $\lambda_{\phi}=0$ (left) and $\lambda_{h\phi}=0$ (right) in Figure 6 shows that achieving a cutoff scale at the GUT scale, $\Lambda_{\text{GUT}}$, requires the DM mass to be approximately $10^{10}$ GeV (left) and $10^{11}$ GeV (right), or larger.
Next, we summarize the more general case in which both the Higgs-portal coupling and the DM four-point self-coupling are present.
Figure 7: Parameter space of $m_{\phi}/m_{h}$ versus $B\Lambda_{\text{GUT}}^{4}$. The left panel compares $B_{\text{grav}}$ (black dashed line) and $B_{\text{non-grav}}$ evaluated with a GUT-scale cutoff (red solid line). The right panel shows the one-loop (black solid line) and two-loop (blue dashed line) contributions to $B_{\text{grav}}$.
Figure 7 illustrates the comparison between the gravitational contribution $B_{\text{grav}}$ (black dashed line) and the non-gravitational contribution $B_{\text{non-grav}}$ evaluated with a GUT-scale cutoff (red solid line) as functions of $m_{\phi}/m_{h}$ in the left panel. The right panel displays the one-loop (black solid line) and two-loop (blue dashed line) contributions to $B_{\text{grav}}$. In both panels, we take $\lambda_{h\phi}=\lambda_{\phi}=1$. Changing the common value of the couplings from unity rescales the vertical axis but does not affect the relative contributions to $B$. From the left panel, we find that the gravitational positivity bound is satisfied for DM masses of order $10^{11}\,\mathrm{GeV}$ or larger when the two couplings are equal. From the right panel, we observe that the two-loop contribution to $B_{\text{grav}}$ becomes dominant for DM masses around $10^{11}\,\mathrm{GeV}$. This behavior is consistent with the case in which only the DM self-coupling is present.
Finally, we consider the case in which the two couplings are varied away from the same value.
Figure 8: Parameter space of $m_{\phi}/m_{h}$ versus $\Lambda$ satisfying the gravitational positivity bound from the $\phi\phi\to\phi\phi$ process. In the left panel, the curves correspond to $\lambda_{\phi}=0$ (red solid), $\lambda_{\phi}=\lambda_{h\phi}/4$ (red dashed), and $\lambda_{\phi}=\lambda_{h\phi}/2$ (red dotted). In the right panel, the curves correspond to $\lambda_{h\phi}=0$ (dark green solid), $\lambda_{h\phi}=\lambda_{\phi}/4$ (dark green dashed), and $\lambda_{h\phi}=\lambda_{\phi}/2$ (dark green dotted).
Figure 8 shows the parameter space of $m_{\phi}/m_{h}$ versus $\Lambda$ satisfying the gravitational positivity bound from the $\phi\phi\to\phi\phi$ process. The left panel demonstrates how increasing the DM self-coupling shifts the minimum DM mass required to realize a GUT-scale cutoff from approximately $10^{10}\,\mathrm{GeV}$ to $10^{11}\,\mathrm{GeV}$. The right panel shows that increasing the Higgs-portal coupling affects the minimum cutoff scale primarily for DM masses near the Higgs boson mass, but does not lead to a dramatic change.
We therefore conclude that, in the presence of both interactions, achieving a cutoff scale at the GUT scale generally requires a DM mass of order $10^{10}$$10^{11}\,\mathrm{GeV}$ (or above), with the larger values favored when the four-point self-coupling plays a significant role.